Wednesday, 19 August

Wednesday, 19 August

Invited Lectures

\underline{14:00 – 14:45}

05 Leonid Polterovich, Tel Aviv University, Tel Aviv, Israel Geometry on the Group of Hamiltonian Diffeomorphisms (I: p.77) Chair: Helmut H.W. Hofer

MA 001

06 Robert A. Oliver, Université Paris-Nord, Villetaneuse, France Vector bundles over classifying spaces (I: p.84) Chair: Charles B. Thomas

MA 004

07 Minoru Wakimoto, Kyushu University, Fukuoka, Japan Representation theory of affine superalgebras at the critical level (I: p.96) Chair: Richard E. Borcherds

MA 005

09 Zhihong Xia, Northwestern University, Evanston, USA Arnold diffusion and variational construction of connecting orbits (I: p.120) Chair: Klaus Kirchgässner

H 1012

11 Gian Michele Graf, ETH Zürich, Zürich, Switzerland Stability of matter in classical and quantized fields (I: p.134) Chair: Volker Bach

H 1028

14 Joan Feigenbaum, AT&T Labs Research, Florham Park, USA Games, Complexity Classes, and Approximation Algorithms (I: p.167) Chair: Hans J. Prömel

H 2032

\underline{15:00 – 15:45}

05 Yongbin Ruan, University of Wisconsin-Madison, Madison, USA Quantum cohomology and its applications (I: p.77) Chair: Helmut H.W. Hofer

MA 001

06 Ronald A. Fintushel, Michigan State University, East Lansing, USA Ronald J. Stern, University of California, Irvine, USA Knot Surgery on 4-Manifolds (I: p.86) Chair: Ulrich Koschorke

MA 004

07 Kari Vilonen, Brandeis University, Waltham, USA Topological Methods in Representation Theory (I: p.95) Chair: Wolfgang Soergel

MA 005

09 Welington de Melo, IMPA, Rio de Janeiro, Brazil Renormalization and Rigidity in one dimensional smooth dynamical systems (I: p.112) Chair: Bernold Fiedler

H 1012

11 Barry McCoy, State University of New York, Stony Brook, USA Rogers-Ramanujan identities: A Century of progress from mathematics to physics (I: p.135) Chair: Robert Schrader

H 1028

14 Toniann Pitassi, University of Arizona, Tucson, USA Combinatorial and algebraic techniques in proof complexity (I: p.171) Chair: Helmut Alt

H 2032

\underline{16:15 – 17:00}

04 & 11 Robbert Dijkgraaf, University of Amsterdam, Amsterdam, The Netherlands The Mathematics of Fivebranes (I: p.62) Chair: Fedor Smirnov MA 005 07 Laurent Lafforgue, Université Paris-Sud, Orsay, France Chtoucas de Drinfeld et applications (I: p.92) Chair: James Arthur

MA 004

08 Michael T. Lacey, Georgia Institute of Technology, Atlanta, USA On the Bilinear Hilbert Transform (I: p.99) Chair: Elias M. Stein

MA 001

10 Fabrice Bethuel, Université Paris-Sud, Orsay, France Recent results for Ginzburg-Landau models (I: p.121) Chair: Dmitri Yafaev

H 1012

12 David John Aldous, University of California, Berkeley, USA Stochastic Coalescence (I: p.140) Chair: Friedrich Götze

H 1028

17 Jorge Nocedal, Northwestern University, Evanston, USA Nonlinear Optimization: The Interplay Between Mathematical Characterizations and Algorithms (I: p.193) Chair: Margaret H. Wright

H 2032

\underline{17:15 – 18:00}

04 Johan de Jong, Princeton University, Princeton, USA Barsotti-Tate groups and crystals (I: p.61) Chair: Gerard van der Geer

MA 005

07 Robert Edward Kottwitz, University of Chicago, Chicago, USA Harmonic analysis on semisimple p-adic Lie algebras (I: p.91) Chair: Michael Rapoport

MA 004

08 Kristian Seip, Norwegian University of Science and Technology, Trondheim, Norway Developments from Nonharmonic Fourier Series (I: p.105) Chair: Klaus D. Bierstedt

MA 001

10 Robert R. Jensen, Loyola University of Chicago, Chicago, USA Viscosity solutions of elliptic partial differential equations (I: p.122) Chair: Stefan Müller

H 1012

12 Jayanta Kumar Ghosh, Indian Statistical Institute, Calcutta, India and Purdue University, West Lafayette, USA Bayesian Density Estimation (I: p.143) Chair: Vladimir I. Koltchinskii

H 1028

17 Jochem Zowe, Universität Erlangen, Erlangen, Germany What is the Best Material? — The Answer of a Mathematician (I: p.196) Chair: Josef Stoer H 2032